OpenStudy (anonymous):

i need the solution set for this radical equation:: sqrt{4x/3}=8

OpenStudy (angela210793):

4x/3=64 4x=64*3=192 x=192/4=48

OpenStudy (anonymous):

and also this one:: 4sqrt{2x-3}=20

OpenStudy (angela210793):

32x-48=400 32x=448 x=448/32 x=14 A={14}

OpenStudy (anonymous):

and sqrt{x/4}=5

OpenStudy (angela210793):

x/4=25 x=100 A={100}

OpenStudy (anonymous):

and 3sqrt{2x-2}=12

OpenStudy (anonymous):

4x/3=64 4x=64*3=192 x=192/4=48

OpenStudy (angela210793):

18x-18=144 18x=144+18=162 x=9 A={9}

OpenStudy (anonymous):

& sqrt{6x-2}=4

OpenStudy (angela210793):

A={3}

OpenStudy (anonymous):

and sqrt{6x+4}=4

OpenStudy (angela210793):

A={2}